If x= cos^2 tetha /(1-cos^2teta) + 2csc^2teta and y = 2cot^2 teta, express y in terms of x
1
Expert's answer
2020-09-20T17:57:17-0400

We'll use the fundamental trigonometric identities:

"sin^2 \theta +cos^2 \theta =1, \textrm{hence } 1-cos^2 \theta = sin^2 \theta."

"1+cot^2 \theta = csc^2 \theta,"

and "cot \theta =\frac{cos \theta}{sin \theta}."

Now we can modify the expression for x:

"x= \frac{cos^2 \theta }{1-cos^2\theta} + 2csc^2\theta=\frac{cos^2 \theta }{sin^2\theta} + 2csc^2\theta=
cot^2 \theta+2(1+cot^2 \theta)="

"=cot^2 \theta+2+2cot^2 \theta=3 cot^2 \theta +2=\frac{3}{2}(2cot^2 \theta)+2=\frac{3}{2}y+2."

Hence

"x=\frac{3}{2}y+2,"

"2x=3y+4,"

"3y=2x-4,"

"y=\frac{2x-4}{3}."


Answer. "y=\frac{2x-4}{3}."